OpenStudy (anonymous):

Help me find the range of each function for the given domain. Please help... I have absolutely no idea what this even is. I give medals.(: f(x) = -2x + 1; {-6, -2, 0} f(x) = x3 + 1; {-2, -1, 3}

OpenStudy (jdoe0001):

f(x) = -2x + 1; {-6, -2, 0} meaning x = -6, x = -2, x = 0 f(-6) = -2(-6)+1 f(-2) = -2(-2)+1 f(0) = -2(0) + 1 and the same for the next function, the values in the braces, are the "x" values, or domain

OpenStudy (anonymous):

Thanks... but I'm still incredibly confused.